OpenStudy (anonymous):

sqrt(72+36e^(12t)+36e^(-12t))=?

OpenStudy (anonymous):

@phi

OpenStudy (phi):

what do you want to do with this? we could re-write it using cosh \[ \cosh(x) = \frac{ e^x + e^{-x}}{2} \] to get \[ \sqrt{72+ 72 \cdot \frac{ e^{12t} + e^{-12t}}{2} } \\ \sqrt{72+ 72 \cosh(12t) } \] there are these identities https://en.wikipedia.org/wiki/Hyperbolic_function#Useful_relations in particular: \[ \cosh(2x) = 2 \cosh^2(x) -1 \] to write \[ \cosh(12t) = 2 \cosh^2(6t) - 1 \] factor out 36 from the square root: \[ 6 \sqrt{2+ 2 \cosh(12t) } \\ 6 \sqrt{2+ 2(2 \cosh^2(6t) - 1)} \\ 6 \sqrt{4\cosh^2(6t)}\\ 12 \cosh(6t) \] if the cosh is positive.